OpenStudy (anonymous):

given secӨ=3, use trigonometric identities to find the exact value of tan^2Ө and sin^2Ө

OpenStudy (cwrw238):

sec^2Ө = 1 + tan^2Ө

OpenStudy (cwrw238):

- will give you value of tan^2Ө

OpenStudy (cwrw238):

sec^2Ө = 1 / cos^2Ө and sin^2Ө = 1 - cos^2Ө will give you the second part

OpenStudy (anonymous):

what did you get?